IF α,B,C are roots of the equation x3-x-1=0
 
FIND THE EQUATION HAVING ROOTS 
(1+α)/(1-α) , (1+B)/(1-B) , (1+C)/(1-C)..
PROVIDE DETAILED SOLUTION IF YOU WANT AN APPROVAL

 

Grade:10

1 Answers

Arun
4 years ago
Dear student
 
now x = 1 + a / 1 – a
 
hence 
x – ax = 1 + a
 
a = x – 1/ (x + 1)
 
hence put (x – 1)/(x +1) in place of x and you will find answer